I have a question that states find the domain and range of the function f(x)=-4x^2+2. How do I do this?
Answer by mala_kumarkos@yahoo.com(10)

f(x)=-4x^2+2
the values of x are called domain
the values of f(x) obtained are called range
if x=(0,1,2,3,4,5,6...)then f(1)=-4(1)+2=-4+2=-2
f(0)=-4(0)+2=0+2 =+2
f(2)=-4(2)+2=-8+2=-6
f(3)=-4(3)+2=-12+2=-10
f(4)=-4(4)+2=-16+2=-14
therefore domain is =(0,1,2,3,4........)
range is =(+2,-2,-6,-10,-14,-18,......)